what is the equation of the line parallel to the line y+1=3(x-5) passing through the point (-2,-5) . . I cannot find any examples in my book on how to solve this.
Answer by mananth(16946)

y+1=3(x-5) passing through the point (-2,-5)
y+1=3(x-5)
y+1=3x-15
-1 -1
y = 3x-15-1
y=3x-16
compare with equation y = mx +b
m=3
..
a parallel line will have same slope.
passing through (-2,-5)
plug the value of x,y,m to find the value of b
-5 = 3*-2+b
-5=-6+b
+6 =6
1=b
..
Now you have slope m=3 and b =1
equation will be
y = 3x+1
...
m.ananth@hotmail.ca